OpenStudy (anonymous):

solve using elimination method 0.3x-0.2y=4 0.5x+0.5y=-15/19

OpenStudy (aravindg):

-1.0048 =x

OpenStudy (aravindg):

y=2.44 approx

OpenStudy (anonymous):

ok that helps. i also have the option to put infinitely many solutions of no solution.

OpenStudy (anonymous):

or*

OpenStudy (saifoo.khan):

\[\Huge x = \frac{820}{19}, y = \frac{-850}{19}\]

OpenStudy (anonymous):

i have to have is as an ordered pair

OpenStudy (anonymous):

thanks you guys!!!! :)

OpenStudy (anonymous):

0.3x - 0.2y=4 multiply by 0.5 --> 0.15 x-0.10y=2.0 (1) 0.5x+0.5y=-15/19 multiply by 0.2 -->0.10x+0.10y =- 0.15 (2) ==>(1)+(2) 0.25x=1.85 x=7.4